# Multimodal Representation Learning for Healthcare

### v1

@@ 1 Introduction
For individuals undergoing cancer systemic therapy (such as chemotherapy), the period between clinical appointments unfolds at home and is often fraught with uncertainty and the risk of unmonitored side effects or rapid health deterioration [oakley2017neutropenic]. In the meantime, due to physical distance, clinicians have limited visibility into their patients’ at-home daily well-being and the subtle onset of treatment-related side effects, creating a critical gap in cancer care [chen2023symptom].
RPM has the potential to address this problem by using multiple sources (such as wearable sensors, mobile apps, and other digital tools) to gather at-home health data [rockey2025effect]. To unlock the true potential of RPM, multi-modal AI was involved to fuse sources together to generate actionable decisions [soenksen2022integrated]. This approach is highly likely to fill the gap because it mirrors the multi-aspect reasoning process of clinicians at home environments.
Previous research in multi-modal AI RPM in cancer care handled the multi-modalities poorly. Some recent research was based on single modality (wearable only [cay_harnessing_2024, low2025consumer], survey only [kairis2025daily, rockey2025effect]). Some research had to reduce sampling frequency because certain modality was sampled slower than the others [jacobsen2023wearable, liu2023evaluating]. There is a lack of standardised framework to address this.
@@ 2 Methods
A prospective observational trial (IRAS 312296) was designed based on the HALO-X multi-modal data collection platform [halox2025infra]. The trial aimed at testing the feasibility of the HALO-X platform in patients with advanced cancer undergoing systemic therapy. Patient recruitment started in March 2023 (inclusion & exclusion criteria see appendix A.1). Detailed collection process is illustrated in Figure 1. Multi-modal data was collected and combined via the platform: baseline characteristics, wearable data in 5-minute epochs, surveys, and clinical events. Surveys used were the 15-item Quality of Recovery questionnaire (QoR-15, 0-10 scale per item, 10=best) [stark2013development] and wellness check-in (an optional yes/no question to indicate overall health today). Treatment log contains information from treatment events such as treatment type (chemotherapy, hormone therapy, immunotherapy etc.), health-caused dose reduction or treatment delay (e.g. low blood platelets).
@@
Data stream asynchronicity is a key consideration when building multi-modal AI for RPM data. The problem appears as: wearable sensors sampled every 5 minutes, survey responses arrived daily, and clinical events occurred irregularly across months of treatments. The way that traditional supervised learning model works either causes the least frequent feature to be re-used numerous times or binning the most frequent feature to a lower frequency (which significantly reduces temporal resolution). Additionally, patients with poor health are known to use monitoring services less [McClaine2024, goldberg2021data, Ortiz2024]. This introduces missingness not at random (MNAR). MNAR in healthcare is notorious for prevalence and tricky to solve [little2002statistical, sterne2009multiple]. To develop an AI model tailored to these data characteristics, we took the concept and adapted a transformer-architecture from the STraTS [tipirneni2022self] with some key adjustments. First, tokenizer encodes each observation as an independent token, creating token sequences where each modality contributes tokens at its native sampling rate, avoiding the first problem mentioned. Second, because each observation is tokenized independently, missing observations can be simply skipped rather than forced to be imputed, avoiding the second problem mentioned. Third, we explicitly engineered missingness features (daily device wearing percentage and continuous absence duration) as additional tokens to enhance missingness learning. The model architecture and training/evaluation pipeline is illustrated in Figure 3.

### v2

@@ 1 Introduction
For individuals undergoing cancer systemic therapy (such as chemotherapy), the period between clinical appointments unfolds at home and is often fraught with uncertainty and the risk of unmonitored side effects or rapid health deterioration [10]. In the meantime, due to physical distance, clinicians have limited visibility into their patients’ at-home daily well-being and the subtle onset of treatment-related side effects, creating a critical gap in cancer care [2].
RPM has the potential to address this problem by using multiple sources (such as wearable sensors, mobile apps, and other digital tools) to gather at-home health data [12]. To unlock the true potential of RPM, multi-modal AI was involved to fuse sources together to generate actionable decisions [13]. This approach is highly likely to fill the gap because it mirrors the multi-aspect reasoning process of clinicians at home environments.
Previous research in multi-modal AI RPM in cancer care handled the multi-modalities poorly. Some recent research was based on single modality (wearable only [1, 8], survey only [5, 12]). Some research had to reduce sampling frequency because certain modality was sampled slower than the others [4, 7]. There is a lack of standardised framework to address this.
@@ 2 Methods
A prospective observational trial (IRAS 312296) was designed based on the HALO-X multi-modal data collection platform [14]. The trial aimed at testing the feasibility of the HALO-X platform in patients with advanced cancer undergoing systemic therapy. Patient recruitment started in March 2023 (inclusion & exclusion criteria see appendix A.1). Detailed collection process is illustrated in Figure 1. Multi-modal data was collected and combined via the platform: baseline characteristics, wearable data in 5-minute epochs, surveys, and clinical events. Surveys used were the 15-item Quality of Recovery questionnaire (QoR-15, 0-10 scale per item, 10=best) [15] and wellness check-in (an optional yes/no question to indicate overall health today). Treatment log contains information from treatment events such as treatment type (chemotherapy, hormone therapy, immunotherapy etc.), health-caused dose reduction or treatment delay (e.g. low blood platelets).
@@
Data stream asynchronicity is a key consideration when building multi-modal AI for RPM data. The problem appears as: wearable sensors sampled every 5 minutes, survey responses arrived daily, and clinical events occurred irregularly across months of treatments. The way that traditional supervised learning model works either causes the least frequent feature to be re-used numerous times or binning the most frequent feature to a lower frequency (which significantly reduces temporal resolution). Additionally, patients with poor health are known to use monitoring services less [9, 3, 11]. This introduces missingness not at random (MNAR). MNAR in healthcare is notorious for prevalence and tricky to solve [6, 16]. To develop an AI model tailored to these data characteristics, we took the concept and adapted a transformer-architecture from the STraTS [17] with some key adjustments. First, tokenizer encodes each observation as an independent token, creating token sequences where each modality contributes tokens at its native sampling rate, avoiding the first problem mentioned. Second, because each observation is tokenized independently, missing observations can be simply skipped rather than forced to be imputed, avoiding the second problem mentioned. Third, we explicitly engineered missingness features (daily device wearing percentage and continuous absence duration) as additional tokens to enhance missingness learning. The model architecture and training/evaluation pipeline is illustrated in Figure 3.
@@ 4 Discussion
We would also like to acknowledge the support of the Urology Foundation, the Champniss Foundation and the Champniss Charitable Trust through the grant, Testing The Relation Between The Host Immune Response And Activity In Patients Receiving Chemotherapy For Solid Cancers.